# A framework for the valuation of insurance liabilities by production cost

### v1

@@ 7 Conclusion
We have introduced a framework for the valuation of insurance liabilities by production cost for an insurance company subject to regulatory solvency capital requirements and insolvency laws. It considers fulfillment only in “sufficiently many cases”, capital, insolvency, and illiquid assets in addition to tradables, in discrete time. The framework is elementary and in particular does not assume the existence of “risk-neutral” pricing measures. It is intended to be practically applicable and, through specifying the production strategies explicitly, allows for a concrete fulfillment by production of the insurance liabilities in practice.
The framework assumes that liability cash flows occur and tradables are traded at discrete dates and that insolvency and regulatory solvency capital requirements are assessed annually. It is based on a fulfillment condition, defining when fulfillment is “sufficient”, and a financiability condition, defining the return requirements of capital investors. Defining the central notion of production strategies in the general case requires distinguishing between positive shares of tradables, which are assets, and short positions of tradables, which are liabilities, and making explicit the assumptions on taking on and closing out short positions. We define production strategies initially for portfolios with positive shares of tradables and extend it to portfolios with non-negative values.
We show that, under suitable assumptions on the financiability condition and the applicable insolvency laws, a production strategy for an arbitrary fulfillment condition can be extended to “almost sure” fulfillment of liabilities that are adjusted only on the complement of the “sufficiently many cases” and in a sense correspond to a “redefinition” of the liabilities that is “allowed” by the given fulfillment condition. We identify the conditions under which the valuation resulting from the production cost can be viewed as an extension of the valuation by market price for investment strategies of tradables. We investigate the implications of the treatment of illiquid assets in the production on the resulting production cost and sketch how the valuation approaches for insurance liabilities from Solvency II and the SST can be seen as special cases of the framework by means of a fulfillment condition that corresponds to the corresponding regulatory solvency capital requirement and a specific “cost-of-capital” fulfillment condition, potentially using additional simplifying assumptions.
@@ 8.1 Production strategies covering failure - definitions, proposition, proof
We assume that according to the applicable insolvency laws, in case of balance sheet insolvency $A^{\prime}_{i}<L_{i}$ at a date $i$, all outstanding incoming and outgoing cash flows of the liabilities $\mathcal{L}$ are proportionally reduced according to Definition 8.1 by essentially the same factor such that balance sheet insolvency is just removed, leading to adjusted liabilities ~$\widetilde{\mathcal{L}}^{(i)}$. The adjustment applies inductively forward in time: after liabilities have been reduced at date $i$, failure may under some developments occur again at later dates. So liabilities ~$\widetilde{\mathcal{L}}^{(i)}$ are successively adjusted ultimately to adjusted liabilities $\widetilde{\mathcal{L}}$. The adjusted liabilities $\widetilde{\mathcal{L}}$ depend on the strategy $\phi$, the illiquid assets $\psi$ and, in our approach, additionally an investment strategy $\theta$, as we explain below. As we show in Proposition 8.2, the adjusted liabilities can be produced by suitably scaling the original production strategy $\phi$. The approach is slightly complicated because the illiquid assets $\psi$ cannot just be scaled down as by assumption they cannot be reduced as they cannot be sold. This requires selecting a strategy $\theta\equiv\theta^{\psi}$ in which the excess $(1-\lambda_{i})\cdot\tilde{Z}^{\psi}_{t}$ of the incoming cash flows $\tilde{Z}^{\psi}_{t}$ above the scaled down cash flows $\lambda_{i}\cdot\tilde{Z}^{\psi}_{t}$ is invested.
@@
where the cash flows $\tilde{Z}^{\theta}_{t}$ are incoming to $\theta$ from $\psi$ and $X^{\theta}_{i}$ are outgoing of $\theta$ but remain within the company’s assets. Such a strategy can always be defined as, at any date $i$, it only pays out its value and the incoming cash flows.
Note that, at date $t=i$, incoming cash flows of the liability $\mathcal{L}$ are scaled by $\lambda_{i-1}$ and outgoing by $\lambda_{i}$, which means that incoming premiums at date $i$ are assumed to be considered in full before failure is declared, whereas outgoing claims, benefits and costs are subsequently reduced, including those at date $i$. For the remaining dates $t\in D\setminus\{0,1,\ldots,T\}$, the same scaling factors are used. In Definition 8.1, liability cash flows are only scaled down over time, although the situation may improve in some developments.
@@
Let the financiability condition be positively homogeneous. Let ϕ+italic-$\phi\in\mathcal{R}^{+}$ with illiquid assets $\psi$ and capital amounts $\mathcal{C}=(C_{i})_{i=i_{min}}^{i_{max}-1}$ be a non-negative production strategy from $i_{min}$ to $i_{max}$ for the liability $\mathcal{L}$ with terminal value $Y_{i_{max}}$. Assume that the production cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ for all $i\in\{i_{min},\ldots,i_{max}-1\}$ are available for all states. Given any fixed non-negative strategy $\theta\equiv\theta^{\psi}$ as in Definition 8.1, define the strategy ϕ~+normal-~italic-$\widetilde{\phi}\in\mathcal{R}^{+}$ for $t\in D[i_{min},i_{max}]$ by
$\widetilde{\phi}_{t}=\lambda_{\lceil t-1\rceil}\cdot\phi_{t}\mbox{ and including }X^{\theta}_{t}\mbox{ as incoming cash flows}$ (32)
with illiquid assets $\widetilde{\psi}_{t}=\lambda_{\lceil t-1\rceil}\cdot\psi_{t}$, capital amounts $\widetilde{\mathcal{C}}=(\lambda_{i}\cdot C_{i})_{i=i_{min}}^{i_{max}-1}$ and terminal value $\tilde{Y}_{i_{max}}=\lambda_{i_{max}}\cdot Y_{i_{max}}$. Then, the strategy ϕ~+normal-~italic-$\widetilde{\phi}\in\mathcal{R}^{+}$ extends $\phi$ to a production strategy for the almost sure fulfillment condition and the same financiability condition for the liabilities ~~$\widetilde{\mathcal{L}}\equiv\widetilde{\mathcal{L}}(\phi,\psi,\mathcal{C},\theta)$ and has production cost
@@
In the proposition, $\widetilde{\phi}_{t}$ being defined as “including $X^{\theta}_{t}$ as incoming cash flows” means that, at a date $i$ ($X^{\theta}_{t}$ is zero at other dates), the cash flows $X^{\theta}_{t}$ flow into the strategy. This does not change the strategy because the additional cash flows are immediately paid out to the capital investors through $C_{i}^{\prime}$. In that sense, the capital investors potentially receive “too much”.
@@
The adjusted liabilities $\widetilde{\mathcal{L}}$ and the strategy $\widetilde{\phi}$ are defined inductively for increasing $i\in\{i_{min}-1\}\cup I$ with $I\equiv\{i_{min},\ldots,i_{max}\}$ by adjusted liabilities ~$\widetilde{\mathcal{L}}^{(i)}$, strategies $\widetilde{\phi}^{(i)}$, illiquid assets $\psi^{(i)}_{t}$, suitable capital amounts $C_{j}^{(i)}$, and strategies $\theta^{(i)}_{t}$, for a suitable terminal value $\tilde{Y}_{i_{max}}^{(i)}$. The quantities in step $i$ account for failure - by adjusting the liabilities such that the almost sure fulfillment condition is satisfied - up to and including date $i$, but potentially not at later dates. This is achieved by scaling all quantities at step $i$ instead of by $\lambda_{j}$ by scaling factors $\lambda_{j}^{(i)}$ for $j\in I$ that are equal to $\lambda_{j}$ for $j\leq i-1$ and equal to $\lambda_{i}$ for $j\geq i$ and with $\lambda_{i_{min}-1}^{(i)}=\lambda_{i_{min}-1}=1=\lambda_{j}^{(i_{min}-1)}$. The strategy $\theta_{t}^{(i)}$ satisfies $\theta_{t}^{(i)}=\theta_{t}$ for $t\leq i+1$ and for the values of the assets and liabilities, $A^{\prime(i)}_{j}=\lambda_{j-1}^{(i)}\cdot A^{\prime}_{j}$ and $L_{j}^{(i)}=\lambda_{j}^{(i)}\cdot L_{j}$ for $j\in I$. In the iteration step from date $i$ to $i+1$, accounting for failure is extended to date $i+1$ without changing the situation up to date $i$. To this end, we define $\xi_{j}^{(i+1)}$ to be $\xi_{j}$ for $j\leq i$,
@@
We show that $\tilde{\phi}^{(i+1)}$ is a production strategy for the liabilities ~$\widetilde{\mathcal{L}}^{(i+1)}$. We have $\tilde{\phi}^{(i+1)}\in\mathcal{R}^{+}$ as it is a non-negative multiple of ϕ+italic-$\phi\in\mathcal{R}^{+}$. To show for $\tilde{\phi}^{(i+1)}$ the equation corresponding to (15) for $t\in D[i_{min},i_{max}]$ with $t>i$ and $t\notin\{i+1,i+2,\ldots\}$, we use $X^{\theta^{(i+1)}}_{t}=0$ and that for $s\in\{t,\gamma(t)\}$ by definition, $\lambda_{\lceil s-1\rceil}^{(i+1)}=\lambda_{\lfloor s\rfloor}^{(i+1)}=\lambda_{i+1}$. So, each of the two sides of (15) for $\tilde{\phi}^{(i+1)}$ is equal to the corresponding side for $\phi$ multiplied by $\lambda_{i+1}$, so (15) for $\tilde{\phi}^{(i+1)}$ follows from (15) for $\phi$.
@@
Now assume that (37) holds at date $j=i+1$. This implies $L_{i+1}^{(i+1)}=\lambda_{i+1}^{(i+1)}\cdot L_{i+1}=\lambda_{i+1}\cdot L_{i+1}=\xi_{i+1}\cdot\lambda_{i}\cdot L_{i+1}=(\lambda_{i}\cdot L_{i+1})\wedge(\lambda_{i}\cdot A_{i+1}^{\prime}+X_{i+1}^{\theta})$. So, $L_{i+1}^{(i+1)}\leq\lambda_{i}\cdot A_{i+1}^{\prime}+X_{i+1}^{\theta}=A_{i+1}^{\prime(i+1)}+X_{i+1}^{\theta^{(i+1)}}$, that is, the almost sure fulfillment condition holds at date $i+1$. We have $C_{i+1}^{\prime(i+1)}\geq(A_{i+1}^{\prime(i+1)}-L_{i+1}^{(i+1)})_{+}=(\lambda_{i}^{(i+1)}\cdot A_{i+1}-\lambda_{i+1}^{(i+1)}\cdot L_{i+1})_{+}$. From $\xi_{i+1}^{(i+1)}\leq 1$ we get $\lambda_{i+1}^{(i+1)}\leq\lambda_{i}^{(i+1)}$, hence $C_{i+1}^{\prime(i+1)}\geq\lambda_{i}^{(i+1)}\cdot(A_{i+1}-L_{i+1})_{+}=\lambda_{i}^{(i+1)}\cdot C_{i+1}^{\prime}$. So we can proceed as above to show that (37) holds at date $i$.

### v2

@@ 6 Conclusion
We have introduced a framework for the valuation of insurance liabilities by production cost for an insurance company subject to regulatory solvency capital requirements and insolvency laws. It considers fulfillment only in “sufficiently many cases”, capital, insolvency, and illiquid assets in addition to tradables, in discrete time. The framework is elementary and in particular does not assume the existence of “risk-neutral” pricing measures. It is intended to be practically applicable and, through explicitly using production strategies, allows for a concrete fulfillment by production of the insurance liabilities in practice.
The framework assumes that liability cash flows occur and tradables are traded at discrete dates and that insolvency and regulatory solvency capital requirements are assessed annually. It is based on two conditions: a fulfillment condition, defining when fulfillment is “sufficient”, and a financiability condition, defining the return requirements of capital investors. Defining the central notion of production strategies in the general case requires distinguishing between positive shares of tradables, which are assets, and short positions of tradables, which are liabilities, and making explicit the assumptions on taking on and closing out short positions. We define production strategies initially for portfolios with positive shares of tradables and extend it to portfolios with non-negative values.
We show that, under suitable assumptions on the financiability condition and the applicable insolvency laws, a production strategy for an arbitrary fulfillment condition can be extended to “full” fulfillment of liabilities that are adjusted only on the complement of the “sufficiently many cases” and in a sense correspond to a “redefinition” of the liabilities that is “allowed” by the given fulfillment condition. We identify the conditions under which the valuation resulting from the production cost can be viewed as an extension of the valuation by market price for investment strategies of tradables. We investigate the implications of the treatment of illiquid assets in the production on the resulting production cost and sketch how the valuation approaches for insurance liabilities from Solvency II and the SST can be seen as special cases of the framework by means of a fulfillment condition that corresponds to the regulatory solvency capital requirement and a specific “cost-of-capital” fulfillment condition, potentially using additional simplifying assumptions.
@@ 7.1 Production strategies covering failure - definitions, proposition, proof
We assume that according to the applicable insolvency laws, in case of balance sheet insolvency $A^{\prime}_{i}<L_{i}$ at a date $i$, all outstanding cash inflows and outflows of the liabilities $\mathcal{L}$ are proportionally reduced according to Definition 7.1 by essentially the same factor such that balance sheet insolvency is just removed, leading to adjusted liabilities ~$\widetilde{\mathcal{L}}^{(i)}$. The adjustment applies inductively forward in time: after liabilities have been reduced at date $i$, failure may under some developments occur again at later dates. So liabilities ~$\widetilde{\mathcal{L}}^{(i)}$ are successively adjusted ultimately to adjusted liabilities $\widetilde{\mathcal{L}}$. The adjusted liabilities $\widetilde{\mathcal{L}}$ depend on the strategy $\phi$, the illiquid assets $\psi$ and, in our approach, additionally an investment strategy $\theta$, as we explain below. As we show in Proposition 7.2, the adjusted liabilities can be produced by suitably scaling the original production strategy $\phi$. The approach is slightly complicated because the illiquid assets $\psi$ cannot just be scaled down as by assumption they cannot be reduced as they cannot be sold. This requires selecting a strategy $\theta\equiv\theta^{\psi}$ in which the excess $(1-\lambda_{i})\cdot\tilde{Z}^{\psi}_{t}$ of the cash inflows $\tilde{Z}^{\psi}_{t}$ above the scaled down cash flows $\lambda_{i}\cdot\tilde{Z}^{\psi}_{t}$ is invested.
@@
where the cash flows $\tilde{Z}^{\theta}_{t}$ are incoming to $\theta$ from $\psi$ and $X^{\theta}_{i}$ are outgoing of $\theta$ but remain within the company’s assets. Such a strategy can always be defined as, at any date $i$, it only pays out its value and the cash inflows.
Note that, at date $t=i$, cash inflows of the liability $\mathcal{L}$ are scaled by $\lambda_{i-1}$ and cash outflows by $\lambda_{i}$, which means that premiums at date $i$ are assumed to be considered in full before failure is declared, whereas outgoing claims, benefits and costs are subsequently reduced, including those at date $i$. For the remaining dates $t\in D\setminus\{0,1,\ldots,T\}$, the same scaling factors are used. In Definition 7.1, liability cash flows are only scaled down over time, although the situation may improve in some developments.
@@
Let the financiability condition be positively homogeneous. Let $\phi\in\mathcal{R}^{\geq 0}$ with illiquid assets $\psi$ and capital amounts $\mathcal{C}=(C_{i})_{i=i_{min}}^{i_{max}-1}$ be a non-negative production strategy from $i_{min}$ to $i_{max}$ for the liability $\mathcal{L}$ with terminal value $Y_{i_{max}}$. Assume that the production cost $\bar{v}_{i}^{\phi,\psi,\mathcal{C}}(\mathcal{L})$ for all $i\in\{i_{min},\ldots,i_{max}-1\}$ are available for all states. Given any fixed non-negative strategy $\theta\equiv\theta^{\psi}$ as in Definition 7.1, define the strategy $\widetilde{\phi}\in\mathcal{R}^{\geq 0}$ for $t\in D[i_{min},i_{max}]$ by
$\widetilde{\phi}_{t}=\lambda_{\lceil t-1\rceil}\cdot\phi_{t}\mbox{ and including }X^{\theta}_{t}\mbox{ as cash inflows}$ (47)
with illiquid assets $\widetilde{\psi}_{t}=\lambda_{\lceil t-1\rceil}\cdot\psi_{t}$, capital amounts $\widetilde{\mathcal{C}}=(\lambda_{i}\cdot C_{i})_{i=i_{min}}^{i_{max}-1}$ and terminal value $\tilde{Y}_{i_{max}}=\lambda_{i_{max}}\cdot Y_{i_{max}}$. Then, the strategy $\widetilde{\phi}\in\mathcal{R}^{\geq 0}$ extends $\phi$ to a production strategy for the full fulfillment condition and the same financiability condition for the liabilities ~~$\widetilde{\mathcal{L}}\equiv\widetilde{\mathcal{L}}(\phi,\psi,\mathcal{C},\theta)$ and has production cost
@@
In the proposition, $\widetilde{\phi}_{t}$ being defined as “including $X^{\theta}_{t}$ as cash inflows” means that, at a date $i$ ($X^{\theta}_{t}$ is zero at other dates), the cash flows $X^{\theta}_{t}$ flow into the strategy. This does not change the strategy because the additional cash flows are immediately paid out to the capital investors through $C_{i}^{\prime}$. In that sense, the capital investors potentially receive “too much”.
@@
The adjusted liabilities $\widetilde{\mathcal{L}}$ and the strategy $\widetilde{\phi}$ are defined inductively for increasing $i\in\{i_{min}-1\}\cup I$ with $I\equiv\{i_{min},\ldots,i_{max}\}$ by adjusted liabilities ~$\widetilde{\mathcal{L}}^{(i)}$, strategies $\widetilde{\phi}^{(i)}$, illiquid assets $\psi^{(i)}_{t}$, suitable capital amounts $C_{j}^{(i)}$, and strategies $\theta^{(i)}_{t}$, for a suitable terminal value $\tilde{Y}_{i_{max}}^{(i)}$. The quantities in step $i$ account for failure - by adjusting the liabilities such that the full fulfillment condition is satisfied - up to and including date $i$, but potentially not at later dates. This is achieved by scaling all quantities at step $i$ instead of by $\lambda_{j}$ by scaling factors $\lambda_{j}^{(i)}$ for $j\in I$ that are equal to $\lambda_{j}$ for $j\leq i-1$ and equal to $\lambda_{i}$ for $j\geq i$ and with $\lambda_{i_{min}-1}^{(i)}=\lambda_{i_{min}-1}=1=\lambda_{j}^{(i_{min}-1)}$. The strategy $\theta_{t}^{(i)}$ satisfies $\theta_{t}^{(i)}=\theta_{t}$ for $t\leq i+1$ and for the values of the assets and liabilities, $A^{\prime(i)}_{j}=\lambda_{j-1}^{(i)}\cdot A^{\prime}_{j}$ and $L_{j}^{(i)}=\lambda_{j}^{(i)}\cdot L_{j}$ for $j\in I$. In the iteration step from date $i$ to $i+1$, accounting for failure is extended to date $i+1$ without changing the situation up to date $i$. To this end, we define $\xi_{j}^{(i+1)}$ to be $\xi_{j}$ for $j\leq i$,
@@
We show that $\tilde{\phi}^{(i+1)}$ is a production strategy for the liabilities ~$\widetilde{\mathcal{L}}^{(i+1)}$. We have $\tilde{\phi}^{(i+1)}\in\mathcal{R}^{\geq 0}$ as it is a non-negative multiple of $\phi\in\mathcal{R}^{\geq 0}$. To show for $\tilde{\phi}^{(i+1)}$ the equation corresponding to (17) for $t\in D[i_{min},i_{max}]$ with $t>i$ and $t\notin\{i+1,i+2,\ldots\}$, we use $X^{\theta^{(i+1)}}_{t}=0$ and that for $s\in\{t,\gamma(t)\}$ by definition, $\lambda_{\lceil s-1\rceil}^{(i+1)}=\lambda_{\lfloor s\rfloor}^{(i+1)}=\lambda_{i+1}$. So, each of the two sides of (17) for $\tilde{\phi}^{(i+1)}$ is equal to the corresponding side for $\phi$ multiplied by $\lambda_{i+1}$, so (17) for $\tilde{\phi}^{(i+1)}$ follows from (17) for $\phi$.
@@
Now assume that (52) holds at date $j=i+1$. This implies $L_{i+1}^{(i+1)}=\lambda_{i+1}^{(i+1)}\cdot L_{i+1}=\lambda_{i+1}\cdot L_{i+1}=\xi_{i+1}\cdot\lambda_{i}\cdot L_{i+1}=(\lambda_{i}\cdot L_{i+1})\wedge(\lambda_{i}\cdot A_{i+1}^{\prime}+X_{i+1}^{\theta})$. So, $L_{i+1}^{(i+1)}\leq\lambda_{i}\cdot A_{i+1}^{\prime}+X_{i+1}^{\theta}=A_{i+1}^{\prime(i+1)}+X_{i+1}^{\theta^{(i+1)}}$, that is, the full fulfillment condition holds at date $i+1$. We have $C_{i+1}^{\prime(i+1)}\geq(A_{i+1}^{\prime(i+1)}-L_{i+1}^{(i+1)})_{+}=(\lambda_{i}^{(i+1)}\cdot A_{i+1}-\lambda_{i+1}^{(i+1)}\cdot L_{i+1})_{+}$. From $\xi_{i+1}^{(i+1)}\leq 1$ we get $\lambda_{i+1}^{(i+1)}\leq\lambda_{i}^{(i+1)}$, hence $C_{i+1}^{\prime(i+1)}\geq\lambda_{i}^{(i+1)}\cdot(A_{i+1}-L_{i+1})_{+}=\lambda_{i}^{(i+1)}\cdot C_{i+1}^{\prime}$. So we can proceed as above to show that (52) holds at date $i$.